# Emerging Platforms Meet Emerging LLMs: A Year-Long Journey of Top-Down Development

### v2

@@
Abstract
Deploying machine learning (ML) on diverse computing platforms is crucial to accelerate and broaden their applications. However, it presents significant software engineering challenges due to the fast evolution of models, especially the recent Large Language Models (LLMs), and the emergence of new computing platforms. Current ML frameworks are primarily engineered for CPU and CUDA platforms, leaving a big gap in enabling emerging ones like Metal, Vulkan, and WebGPU.
While a traditional bottom-up development pipeline fails to close the gap timely, we introduce TapML, a top-down approach and tooling designed to streamline the deployment of ML systems on diverse platforms, optimized for developer productivity. Unlike traditional bottom-up methods, which involve extensive manual testing and debugging, TapML automates unit testing through test carving and adopts a migration-based strategy for gradually offloading model computations from mature source platforms to emerging target platforms. By leveraging realistic inputs and remote connections for gradual target offloading, TapML accelerates the validation and minimizes debugging scopes, significantly optimizing development efforts.
TapML was developed and applied through a year-long, real-world effort that successfully deployed significant emerging models and platforms. Through serious deployments of 82 emerging models in 17 distinct architectures across 5 emerging platforms, we showcase the effectiveness of TapML in enhancing developer productivity while ensuring model reliability and efficiency. Furthermore, we summarize comprehensive case studies from our real-world development, offering best practices for developing emerging ML systems.
Index Terms:
@@ Abstract
$\bigstar$\$$\bigstar$\$footnotetext: Equal contribution.
I Introduction
Machine learning (ML) has been ever-evolving to advance the performance of various data-driven applications, from the early days of simple linear classifiers to the latest advancements in large language models [1]. To power large-scale ML, the silicon industry has also been continually expanding the boundaries of what is computationally possible, by innovating the computing platforms for CPUs, GPUs, and emerging domain-specific accelerators [2, 3].
Because there is a huge gap in mapping high-level ML mathematics to low-level machine executables, substantial efforts have been made to develop ML optimizers [4, 5, 6] to efficiently run ML models. These frameworks, given a trained model, deploy it over inference-optimized platforms by providing efficient native kernels (e.g., CUDA/C functions) to compute the operators/sub-graphs of the model. These kernel functions can be either manual-crafted in a native programming language (e.g., CUDA) or compiler-generated from intermediate representations [7, 8, 9, 10] (IRs). Yet, existing ML frameworks have been primarily implemented and optimized for the CPU and CUDA platforms, leaving a blank space for other emerging platforms.
With recent advances and interest in Large Language Models (LLMs), powerful new models of various architectures are released rapidly by large organizations (Figure 1). Meanwhile, an array of emerging computing platforms have been proposed recently to unlock the possibility of running ML models on environments without NVIDIA GPUs, such as Metal for Apple Silicon [11], WebGPU for web browsers [12], etc. It is crucial to enable sustainability by enabling a broader set of emerging platforms both locally and on the server. Emerging local ML provides a fundamental guarantee for user privacy and resolves the cost and network requirements of a server platform. Meanwhile, expanding server-grade computing options for GPU customers offers greater diversity in terms of performance, cost-effectiveness, flexibility, etc.
Engineering an ML framework is hard, and engineering an ML framework to enable emerging ML models on emerging platforms is an even more challenging software engineering problem. The immediate but unavoidable challenge is to learn the new programming primitives from the target platform. Furthermore, software developers can spend 35-50% of their time testing and debugging software [13]. This is also naturally true here as two main challenges arise for testing and debugging. First, sufficient and diverse test cases are required to test the implementation of individual operators, yet oftentimes developers have to craft these tests manually and the whole process is tedious and laborious. Second, the tooling support and ecosystems for emerging platforms are extremely limited (e.g., no debuggers), leading to a painful debugging and analysis experience.
Figure 1: Timeline of selected new model supports in TapML [^2] For clarify we only visualize a subset of important models we supported..
Figure 2: Bottom-up v.s. Top-down approaches to emerging ML system development
Bottom-up ML deployment. Existing ML frameworks have been commonly practicing a bottom-up development style, which unfortunately no longer meets the productivity requirement when supporting emerging ML systems. The top half in Figure 2 illustrates the key steps for bottom-up ML deployment. There are two steps to support a new model in a bottom-up fashion: (i) Developers implement previously unsupported operators; and (ii) implemented operators are put together into a whole model for end-to-end testing and debugging. Specifically, in step one, developers analyze operators to support, implement them, and manually craft tests for validation. Because manually crafted tests often use synthetic inputs whose distribution can substantially differ from those inherited from realistic inputs, operator-wise validation is inadequate. End-to-end model-wise testing is still effective in manifesting obscure bugs such as precision errors; however, bugs detected by model-wise testing are hard to debug, as the symptoms are simple pass-or-fail signals in terms of result consistency, missing details to infer the error-inducing components. As such, troubleshooting the large model space leads to challenging debugging experiences and significantly slows down the development process.
@@ I Introduction
Top-down ML deployment. To this end, we propose TapML, a top-down approach and tooling to enable emerging ML systems, optimized for developer productivity. The bottom half of Figure 2 illustrates the workflow of TapML. At the high level, TapML regards the deployment as a migration task to gradually offload the model implementation from a mature source platform (e.g., CUDA) to an emerging target platform (e.g., WebGPU). Specifically, TapML takes the following steps to deploy an ML model for an emerging platform. First, TapML automates the curation of unit tests through test carving [14], which automatically tracks intermediate computation snapshots of the source-platform execution as test cases for the target-platform implementation. Notably, the inputs to bootstrap the model during instrumentation are realistic inputs (e.g., meaningful sentences for LLMs) such that the precision regression can focus on real-world data distribution. Next, we perform gradual target offloading. We start with a compute graph fully implemented on the source platform which can correctly pass all system tests. Then we gradually offload each operator in the compute graph from the source platform to the target platform, until the target platform fully runs the compute graph. For each operator, we first use the carved unit tests to perform operator-wise validation, after which we substitute the source-platform implementation with the target-platform implementation. Such a flow also provides an error isolation mechanism where a model-wise failure must be introduced by the operator newly implemented on the target platform. As such, developers can quickly localize the error in a small scope and use immediate operator- and model-wise validation feedback for efficient debugging. Meanwhile, to serve the hybrid graph and perform differential testing [15], TapML builds a universal runtime that unifies cross-platform data communication and function calling. The runtime allows us to invoke remote functions on the device and transfer its inputs and outputs between source and target platforms seamlessly in Python as if running a local program. Lastly, the workflow finalizes when the compute graph is fully implemented on the target platform and passes model-wise validation. As is shown in Figure 1, we have been applying TapML starting from 2024 Q1 and have productively supported various emerging ML models on major emerging platforms, unlocking powerful LLM access to users of various devices, even within the same day of model release.
This paper summarizes our lessons in building and deploying real-world open-source ML systems for emerging platforms. Below we summarize our contributions:
@@
  • Dimension: While the optimization and abstraction for ML computation have been well studied in the past decade, we look at the productivity challenge for deploying emerging ML and address this challenge by advancing the software development methodology.
@@
  • Study: We turn our year-long experience of deploying 82 emerging models on 5 novel platforms into summarized case studies, to facilitate the future development of emerging ML systems.
II Challenges
In this section, we demonstrate key challenges and pain points in developing emerging ML systems.
II-A Insufficient and Laborious Testing
Figure 3: Sample unit test in TensorFlow.js.
Developers invest significant effort in writing extensive test suites to ensure the reliability of machine learning (ML) frameworks. For example, 36% source code of PyTorch [5], a well-known ML framework, and 42% source code of TVM [6], a well-known ML compiler, are testing code, as of Q1 2023. As an instance of emerging platforms, TensorFlow.js [16] supports 85 common operators for the WebGPU platform using 18,001 lines of source code. Despite that not all operators but a few important ones (e.g., matmul and conv2d) are unit-tested, 3,521 lines of the source code are making unit tests.
Figure 3 exemplifies a test case in TensorFlow.js which validates the functionality of conv2d. The test case is crafted by developers by specifying operator attributes (Line 2-3), the input data (Line 4-7), and the oracle including the expected output shape (Line 11) and output values (Line 12-14). Yet, this test only covers one path of the operator whereas tons of other tests use different input and operator configurations to exercise different paths. Meanwhile, the manually specified inputs are rather synthetic given the apparent value orders in x and w, which may not accurately reflect realistic data distribution for a specific model, introducing a distribution gap between test cases and real-world scenarios.
This example highlights two pain points in test development: (i) the labor-intensive nature of manual testing; and (ii) the gap between synthetic and realistic inputs. In addition to the number of tests desired, crafting comprehensive test suites requires meticulous consideration of detailed input-output shapes and values, often involving complex operator constraints [17, 18]. Furthermore, relying solely on synthetic inputs for operator validation may not accurately assess the performance of operators with real-world data, especially concerning factors like precision. This gap is particularly concerning when applying quantization techniques [19, 20] to optimize model efficiency at the cost of robustness [21].
To conclude, addressing the challenges in test development for emerging ML systems is crucial to enhance developer productivity and ensure model reliability. Efforts to bridge the gap between synthetic and realistic inputs, alongside automating test development, are essential for the continued advancement and adoption of ML models and platforms.
@@ II-A Insufficient and Laborious Testing
Figure 4: The computation representations in the TapML workflow. To deploy a trained model, TapML converts it into a general compute graph, where each operator can be offloaded to various backends through target code generation.
II-B Obscure End-to-End Bugs
Unit tests do not universally ensure the accuracy of end-to-end models due to the inability to guarantee numerical precision for every possible input. This limitation is exemplified by an observed discrepancy on Qualcomm Hexagon backends: when performing the calculation $1035-1031$ with the float16 data type, the expected result of 4 was erroneously computed as 5. Although this deviation, less than 0.1%, may seem reasonable, it diverges from the behaviors observed on other computational platforms and can have significant implications for the integrity of end-to-end models. It is important to note that such errors are input-specific and not attributable to faults in the kernel implementations themselves, rendering them elusive to detection via operator-level testing.
Furthermore, as ML models increase in complexity and size, incorporating a greater number of operators, the cumulative impact of precision errors becomes more pronounced. Even if an individual operator’s error falls within the tolerable margin of 0.1% at the operator level, the aggregate effect across hundreds of operators can amplify this error to 1% or even 5%. Such levels of inaccuracy are untenable, highlighting the need for rigorous testing protocols that account for the variances in floating-point computations attributable to different hardware and software configurations.
II-C Limited Tooling in Arising Systems
The infrastructure stack in mature platforms has been well-engineered which enables developers to easily invoke device functions, print remote data, transfer data between host and device, and so on. Moreover, mature platforms even develop advanced debugging utilities. For example, Polygraphy [22] is a graph reducer in TensorRT [4]. When a model fails to compile by TensorRT, Polygraphy automatically reduces the full model to a small sub-graph that can reproduce the error, helping developers to quickly localize the fault.
In contrast, because emerging platforms are newly established, the tooling in their ecosystem is rather limited. Consequently, oftentimes it is challenging for developers to reason about internal states of computation performed in such arising platforms, which is important for debugging. Such a fact also deteriorates the prior two challenges, solving which requires advanced tooling to communicate with the underlying platform. Therefore, it is crucial to build ideally a universal middle-end tooling that allows developers to easily inspect and analyze the model execution.
III Design of TapML
In this section, we detail the design of TapML.
III-A Overview
@@
The goal of deploying a model on a specific platform is to describe the model’s computation using the native program of the target platform. Specifically, Figure 4 illustrates a typical deployment workflow in TapML. The input to the deployment flow is a trained model produced by the training framework (e.g., PyTorch). Subsequently, the trained model is translated by the framework into an internal intermediate representation (IR) that generally describes the semantics of the model. Through a series of general-purpose optimizations, the IR flows to the target code generation stage which transforms the computation to native programs of the target platform. For example, existing frameworks often specialize the code generation pipeline to CUDA, a mature platform in the NVIDIA ecosystem, and these computations are finally represented by CUDA C functions, shown in the purple box in Figure 4.
@@
We describe the goal and overall steps of TapML using the example of Figure 4. The goal of TapML is to facilitate the process of implementing and testing target code generation of the model operators on the emerging platform, i.e., on WebGPU such that users can run their model on a browser, as is shown in the blue box in Figure 4. To make this happen, in Section III-B TapML makes basic assumptions for the deployment framework, including a compute graph abstraction and a reference implementation from a mature platform, i.e., CUDA here. To make testing easy, in Section III-C, TapML automatically creates realistic operator-level tests through test carving, which instruments the execution of each operator on CUDA and collects its input-output pairs as unit tests for the WebGPU support. Specifically, TapML views the deployment process as a migration task — in the very beginning we have a reliable CUDA implementation (i.e., pure CUDA compute graph) of the model and we want to eventually run the model fully on WebGPU (i.e., pure WebGPU compute graph). As such, for simplify the mindset, each time we focus on implementing one operator for WebGPU and validate the operator implementation using the automated tests. Furthermore, for correctness, we also want to make sure after putting the newly implemented operator everything still works as well as previously. Therefore, in Section III-D, we maintain an initial CUDA compute graph whose correpsonding operators are gradually offloaded to WebGPU once a new WebGPU operator is implemented, and thus we perform model-wise validation in each step. The gradual offloading mechanism allows developers to focus on a minimal unit (i.e., one operator) each time, which minimizes the debugging scope, and once it passes model-wise validation we are confident that it will not bring hidden bugs to further steps. In Section III-E, TapML also simplifies these procedures through its universal runtime for flexible cross-platform interaction. The universal tooling stack in TapML can easily run remote computation as if locally and exchange useful information (e.g., debugging message) with the target platform.
III-B Prerequisites

### v3

# Productively Deploying Emerging Models on Emerging Platforms: A Top-Down Approach for Testing and Debugging

@@
Abstract.
While existing machine learning (ML) frameworks focus on established platforms, like running CUDA on server-grade GPUs, there have been growing demands to enable emerging AI applications in a broader set of scenarios, such as running Large Language Models (LLMs) within browsers and mobile phones. However, deploying emerging models on new platforms (such as Metal and WebGPU) presents significant software engineering challenges due to rapid model evolution and limited tooling and practices for these platforms.
Previous practice for ML model deployment often follows a bottom-up fashion, where engineers first implement individual required operators and then put them together. However, this traditional development approach fails to meet the productivity requirements when deploying emerging ML applications, with the testing and debugging part as a bottleneck. To this end, we introduce TapML, a top-down approach designed to streamline model deployment on diverse platforms. While the traditional bottom-up approach requires crafting manual tests, TapML automatically creates high-quality, realistic test data through operator-wise test carving. Furthermore, TapML uses a migration-based strategy to gradually offload model implementation from the mature source platform to the target platform, minimizing the debugging scope of compound errors.
TapML has been used as the default development method in the MLC-LLM project to deploy emerging ML models. In the past two years, TapML has accelerated the deployment of 105 emerging models in 27 model architectures across 5 emerging platforms. We show that TapML effectively boosts developer productivity while ensuring the quality of deployed models. Furthermore, we summarize comprehensive case studies from our real-world development, offering best practices for developing emerging ML systems.
@@ Abstract.
1. Introduction
With the recent interest in generative AI, such as Large Language Models (LLMs), major organizations have been releasing and deploying various powerful models. These emerging models have been empowering various emerging applications such as art generation, coding assistance, and autonomous agents for computer use (Anthropic, 2024). Meanwhile, these models of billions of parameters are computationally expensive yet can be served under various computing scenarios. As such, the silicon industry has been continually expanding the boundaries of what is computationally possible, by innovating their computing platforms based on efficient accelerators such as GPU, TPU (Jouppi et al., 2023), and LPU (Abts et al., 2022).
From a software development perspective, deploying machine learning (ML) models is engineering-intensive. Because there is a huge gap in mapping high-level ML mathematics to low-level machine executables, substantial efforts have been made to develop optimized inference engines (NVIDIA, 2022; Paszke et al., 2019; Chen et al., 2018) to serve ML models efficiently. Specifically, these frameworks deploy a given trained model by overloading its computation (e.g., operators and sub-graphs) using platform-optimized native kernel [^1] Kernels in our paper refers to low-level functions that implement the computation of an operator or fused-computation of a sub-group of operators (e.g., implementing convolution plus ReLU in one function). For clarity, the subsequent discussion assumes one kernel corresponds to one operator and uses both terms interchangeably. functions (e.g., implemented in CUDA/C). Such kernel functions can be either manually crafted in a platform-native programming language (e.g., CUDA) or automatically generated by ML compilers (Lattner et al., 2020; Tillet et al., 2019; Chen et al., 2018). Crafting new operator kernels or building a complicated compiler to generate kernels requires extensive efforts from domain experts. Therefore, existing ML frameworks focus on supporting and optimizing a limited set of mainstream platforms such as the CPU and CUDA platforms, leaving other emerging platforms under-supported.
The models behind these emerging AI applications do not only run on CUDA functions and NVIDIA servers. Expanding the computing diversity in serving ML models can unlock a broader range of applications and features. For example, serving LLMs within browsers via the WebGPU (Consortium, 2024) platform allows for local web agents without compromising data privacy. Likewise, the new “Siri” based on Apple Intelligence (Inc., 2024) depends on the neural processing unit to drive energy-efficient on-device machine learning. These examples highlight the growing trend of diversifying computation platforms to support next-generation models and applications.
Deploying ML models is hard, and deploying emerging ML models on emerging platforms presents an even greater software development challenge. Beyond implementing the corresponding operators, significant time and energy are required for testing and debugging both individual operators and the whole end-to-end model. To this end, our work focuses on studying and enhancing the productivity of workflows that support these emerging models. We begin by explaining the current model deployment practices, followed by a demonstration of our solution.
Figure 1. Bottom-up v.s. Top-down approaches to emerging ML system development
Current practice: bottom-up deployment. The existing practice of model deployment naturally follows a bottom-up pipeline, shown in the top half of Figure 1: (i) “Bottom”: Developers analyze and implement the missing operator implementations that are required to serve the under-deployment model on the target platform efficiently; and (ii) “Up”: developers attempt to serve the model using the target-platform operator implementations and test it using end-to-end samples and benchmarks. However, such a strategy presents challenges in both testing and debugging. In the “bottom” phase, developers typically construct synthetic inputs, resulting in test insufficiency in both quality and quantity, due to the synthetic and unscalable nature of manual testing. Meanwhile, deferring model-wise testing after kernel-wise testing can lead to debugging disasters: When numerical inaccuracy occurs in the integral model testing, fault localization becomes challenging. The error could come from inaccuracy caused by faulty single-operator implementations or propagated precision errors due to the co-functioning of multiple operators. As such, troubleshooting in the large operator space can bring significant mental burdens, slowing down the development process.
We argue that productive deployments of emerging ML require innovations in software methodologies and toolings. As such, we introduce TapML, a top-down approach for testing and debugging the deployment of emerging ML systems.
@@ 1. Introduction
Our top-down & incremental practice. The bottom half of Figure 1 illustrates the workflow of TapML. First, we perform test carving on the model’s source platform (i.e., already supported mature platforms such as CUDA) by running the model with real-world inputs (e.g., meaningful sentences and images). During the model execution, we automatically instrument intermediate inputs/outputs as test cases to be used for testing operator implementations on the target platform (i.e., the emerging platform to deploy the model). Next, developers support the under-deployed model on the target platform by regarding it as a platform migration task. The migration process starts by serving the compute graph of the model entirely on its source platform (e.g., CUDA). Each time, we implement one operator for the target platform, test it, and update the hybrid compute graph, until the whole graph entirely runs on the target platform. Specifically, we test each target-platform operator in two steps: (i) test the standalone operator implementation using carved test cases; and (ii) test the end-to-end model via an operator offloading mechanism, i.e., replacing the corresponding source-platform implementation with target-platform implementation in the flow of computation. Such a flow provides an error isolation mechanism where a test failure must be introduced by the operator newly implemented on the target platform. As such, developers can quickly localize the error in a small scope and eagerly debug it while the development context is still fresh. Additionally, emerging platforms typically lack robust developer utilities to support our methodology. Therefore, tooling-wise TapML builds a universal runtime that unifies cross-platform data communication and function calling, which can support hybrid computation graphs during migration and simplifies the testing and debugging on the emerging platform.
We summarize our main contributions as follows:
@@
  • Dimension: While the optimization and abstraction for ML computation have been well studied in the past decade, we look at the productivity challenge for deploying emerging ML and address this challenge by advancing the software development approach.
@@
  • Implementation: We implemented a universal runtime to facilitate the practice of the TapML methodology. Our universal runtime currently supports 5 platforms.
  • Study: TapML has been used as the testing and debugging method behind the MLC-LLM project (MLC team, 2025). Within two years of efforts, we have been applying TapML to deploy 105 emerging models on 5 novel platforms. Based on these experiences, we case study the use of TapML and interesting troubleshooting experiences to facilitate the future development of emerging ML systems.
2. Challenges
In this section, we discuss the main challenges in deploying emerging ML systems and why a bottom-up development methodology can be largely optimized.
2.1. Motivating Example
Consider the deployment of a Llama model (Touvron et al., 2023) in a browser environment through the WebGPU platform. ML engineers typically approach this deployment in a bottom-up fashion:
  1. (1)
    Operator analysis: The very first step is to analyze the model’s operator structure. Models like Llama include complex operators such as embeddings, multi-head attention, root-mean-square layer normalization, and matrix multiplications. In reality, the required implementations can become intricate due to efficiency demands; for instance, matrix multiplication alone may require dozens of specialized variants tailored to scenarios like query-key-value (QKV) and feedforward layers.
  2. (2)
    Operator implementation: Once the necessary operators are identified, engineers implement their computational logic on the WebGPU platform using WebGPU Shading Language (WGSL). The computation logic is often inferred from academic papers, online documents, and existing implementations from other platforms.
  3. (3)
    Operator-wise testing & debugging: Each operator implemented in WGSL undergoes targeted testing. Engineers craft test cases according to the operator’s semantics and validate the operator’s accuracy by running these tests across browser platforms with WebGPU support (e.g., Chrome and Safari). Per the testing results, they debug the operator implementation until all tests pass before moving to other operators.
  4. (4)
    Model-wise testing: After implementing all operators in WGSL, engineers attempt to reproduce full model execution in the browser by scheduling these WGSL functions according to the model’s data flow. Specifically, for language models like Llama, developers test them by gathering prompts (e.g., from open datasets or developer customization) and observing the output coherence and corresponding metrics. Sometimes, even when individual operators pass, model-wise validation may reveal issues [^2] For example, real-world inputs may easily trigger activations (e.g., ReLU) that manual inputs cannot trigger. . Common model-wise errors include faulty operator implementation due to insufficient operator-wise testing, accumulated inaccuracies when propagating data across multiple operators, and even platform bugs (e.g., miscompilation or runtime errors), making debugging especially challenging.
In the follow-up subsections, we discuss the inherent limitations of the traditional bottom-up development pipeline.
2.2. Test Insufficiency
Figure 2. Sample unit test in TensorFlow.js.
Complex systems such as ML frameworks require extensive testing to ensure their reliability. For example, 40-42% of the Python and C++ source code of PyTorch (Paszke et al., 2019) and TVM (Chen et al., 2018) are testing code, as of Q3 2024. While ML frameworks developed comprehensive tests to exercise implementations on mature platforms, setting up emerging models on emerging platforms often suffers from low test coverage due to limited tests and constrained timelines. Meanwhile, manually curating high-quality test cases is labor-intensive for developers. For instance, the WebGPU backend of TensorFlow.js (Google, 2024) is implemented in 19,530 lines of code, with only 3,549 lines of code attributed to testing. Specifically, most testing codes are dominated by a few common operators such as matrix multiplication and convolution. Furthermore, directly migrating tests from one platform to another is also challenging due to inconsistency in API granularity, interface compatibility, and supported data types.
Figure 2 exemplifies a test case for conv2d in TensorFlow.js. Developers craft the test case by specifying operator attributes (Lines 2-3), input data (Lines 4-7), and the oracle, including the expected output shape (Line 11) and output values (Lines 12-13). It is worth noting from Lines 4-7 that the test inputs are rather synthetic, given the apparent value orders in x and w. Such simplistic test inputs may not represent the distribution of inputs derived from real-world user-given data, whose precision deserves more attention.
2.3. Obscure Precision Bugs
Result inconsistency is one of the most challenging and harmful bug symptoms in software engineering. Compared to other fail-stop bugs such as program crashes, debugging result inconsistency is extremely hard as it only presents the error in the final result without associating the intermediate step that introduces the error. Unfortunately, such types of errors can be common in ML due to its heavy use of arithmetic computation (i.e., few operators can trigger crashes) and floating-point numbers. For example, we found the calculation $1035-1031$ in float16 erroneously leads to a result of 5 on the Qualcomm Hexagon backend.
Furthermore, despite passing operator-wise testing, result inconsistency bugs can still happen and can be even more challenging to debug in the end-to-end model execution. This is because emerging models have increasing complexity and size and incorporate a greater number of operators of various data types. Therefore, the cumulative impact of precision errors becomes more pronounced, and a large debugging scope needs to be checked in order to isolate the error.
Meanwhile, in the bottom-up development pipeline, operator-wise testing and model-wise testing are entirely separated into different phases. As such, when debugging model-wise issues, it is challenging for developers to recall related operator-wise debugging contexts. This limitation calls for applying model-wise testing while developing single operators.
2.4. Limited Tooling in Arising Systems
The infrastructure stack in mature platforms has been long developed, allowing developers to seamlessly invoke device functions, transfer data between host and device, and so on. In addition, these platforms can also provide ready-to-use debugging tools. For example, the TensorRT (NVIDIA, 2022) ecosystem implements Polygraphy (NVIDIA, [n. d.]), a graph reducer to localize error-inducing sub-graphs.
In contrast, emerging platforms often have limited tooling support, making it challenging for developers to analyze internal computational states—a crucial aspect of debugging. This lack of tools exacerbates existing challenges. Therefore, developing infrastructures and tooling that facilitate developer interaction with these platforms is essential.
@@ 2.4. Limited Tooling in Arising Systems
Figure 3. The computation representations in the TapML workflow. To deploy a trained model, TapML converts it into a general compute graph, where each operator can be offloaded to various backends through target code generation.
3. Design of TapML
3.1. Overview
@@
The goal of deploying a model on a specific platform is to describe the model’s computation using the native program of the target platform. Specifically, Figure 3 illustrates a typical deployment workflow in TapML. The input to the deployment flow is a trained model produced by the training framework (e.g., PyTorch). Subsequently, the trained model is translated by the framework into an internal intermediate representation (IR) that generally describes the semantics of the model. Through a series of general-purpose optimizations, the IR flows to the target code generation stage, which transforms the computation to native programs of the target platform. For example, existing frameworks often specialize the code generation pipeline to CUDA, a mature platform in the NVIDIA ecosystem, and these computations are finally represented by CUDA C functions, shown in the purple box in Figure 3.
@@
We describe the goal and overall steps of TapML using the example of Figure 3. The goal of TapML is to facilitate the process of implementing and testing target code generation of the model operators on the emerging platform, i.e., on WebGPU such that users can run their model on a browser, as is shown in the blue box in Figure 3. To make this happen, in Section 3.2 TapML makes basic assumptions for the deployment framework, including a compute graph abstraction and a reference implementation from a mature platform, i.e., CUDA here. To make testing easy, in Section 3.3, TapML automatically creates realistic operator-level tests through test carving, which instruments the execution of each operator on CUDA and collects its input-output pairs as unit tests for the WebGPU support. Specifically, TapML views the deployment process as a migration task: in the very beginning, we have a reliable CUDA implementation (i.e., pure CUDA compute graph) of the model, and we want to eventually run the model fully on WebGPU (i.e., pure WebGPU compute graph). As such, to simplify the mindset, each time we focus on implementing one operator for WebGPU and validate the operator implementation using the automated tests. Furthermore, for correctness, we also want to make sure that after putting in the newly implemented operator, everything still works as previously. Therefore, in Section 3.4, we maintain an initial CUDA compute graph whose corresponding operators are gradually offloaded to WebGPU once a new WebGPU operator is implemented, and thus, we perform model-wise validation in each step. The gradual offloading mechanism allows developers to focus on a minimal unit (i.e., one operator) each time, which minimizes the debugging scope, and once it passes model-wise validation, we are confident that it will not bring hidden bugs to further steps. In Section 3.5, TapML also simplifies these procedures through its universal runtime for flexible cross-platform interaction. The universal tooling stack in TapML can easily run remote computation as if locally and exchange useful information (e.g., debugging message) with the target platform.
3.2. Prerequisites